A high energy proton from the sun enters the Earths magnetic field at a speed of 6.0x10^6 m/s perpendicular to the field lines. The field is uniform and with strength B=4.5x10^-5 T.
a) What is the magnetic force on the proton when it first enters the B-field?
b) Calculate the radius of the protons path in the B field.
c) determine the frequency and the period of the protons orbit.
A high energy proton from the sun enters the Earth
Answer #1

a). Since velcoity of proton and Magnetic Field are perpendicular to each other, then

F= q(\vec{v} \times \vec{B})= qvBsin90^{\circ}= qvB

F= (1.6 \times 10^{-19})(6 \times 10^6)(4.5 \times 10^{-5})

  F= 43.2 \times 10^{-18} N

b), Now since the proton will move in a circular path then above magnetic force will be equal to centripetal force acting on it. Thus

qvB= mv^2/r

r= mv/qB

r= (1.67 \times 10^{-27})(6 \times 10^6)/(1.6 \times 10^{-19})(4.5 \times 10^{-5})

  r= 1.391 \times 10^{3} m

c). Also then Time Period of revolution, T= 2\pi r/v

  T= 2(3.14) (1.391 \times 10^3)/(6 \times 10^6)

  T= 1.455 \times 10^{-3} s

and frequency of revolution is f=1/T= 1/(1.455 \times 10^{-3})= 687.28 s^{-1}